OpenStudy (anonymous):

6. In a 45°- 45°- 90° right triangle, the length of the hypotenuse is 15. How long are the legs? 6. In a 45°- 45°- 90° right triangle, the length of the hypotenuse is 15. How long are the legs? @Mathematics

OpenStudy (anonymous):

15/sqrt(2)

OpenStudy (hoblos):

10.6

OpenStudy (anonymous):

\[15\div \sqrt{2}?\]

OpenStudy (anonymous):

huh??? sorry im confused

OpenStudy (hoblos):

it is \[\sqrt{15^{2}\div2}\]

OpenStudy (anonymous):

its a special case of triangle.

OpenStudy (anonymous):

oh ok why" and how do i get the answer

OpenStudy (anonymous):



OpenStudy (anonymous):

for a 45-45-90 triangle the legs = x then the hyptenous = x*sqrt(2)

OpenStudy (anonymous):

so if you know the hypt. of a 45.45.90 then to find the lengths of the legs, divide hypt. by sqrt(2)

OpenStudy (anonymous):

and that gives us radical15^2/2

OpenStudy (anonymous):

that gives us what?

OpenStudy (anonymous):

\[\sqrt{15^{2}\div}2\]

OpenStudy (anonymous):

\[15/\sqrt{2} = 15\sqrt{2}/2\]

OpenStudy (anonymous):

so theres no work for it? if there is can u show me the steps please im sorry but math is so hard for me to understand

OpenStudy (anonymous):

please :'(

OpenStudy (anonymous):

\[15^2=x^2+x^2\]\[15^2 = 2 x^2 \]\[\frac{15^2}{2}=x^2 \]\[\sqrt{\frac{15^2}{2}}=\sqrt{x^2} \]\[\frac{15}{\sqrt{2}}=x\]